# Catalog-Native LLM: Speaking Item-ID dialect with Less Entanglement for Recommendation

### v1

@@ A.3 Generative and Hybrid Recommender Models
Generative recommenders recast recommendation as a sequence generation task (Yang et al., 2025), unifying retrieval and ranking in one model. HSTU (Zhai et al., 2024) employs a Transformer-based transducer, scaling up to 1.5T parameters and achieving large offline and online gains, while demonstrating NLP-like scaling laws for recommendation. TIGER (Rajput et al., 2023) compresses item vocabularies via multi-code vector quantization. OneRec (Deng et al., 2025) unifies retrieval and ranking in an encoder–decoder Transformer with sparse Mixture-of-Experts (MoE) Shazeer et al. (2017); Fedus et al. (2022); Ma et al. (2018); Tang et al. (2020); Xu et al. (2024); Zhang et al. (2024); Wang et al. (2024b) for capacity scaling and adds Iterative Preference Optimization for alignment. These approaches offer novelty, explainability, and unified modeling, but require heavy compute and careful strategies to retain collaborative memory.
@@ B.1 Baselines
Early sequential modeling. GRU4Rec (Hidasi et al., 2015) pioneers GRU-based session modeling; SASRec (Kang & McAuley, 2018) introduces unidirectional self-attention; BERT4Rec (Sun et al., 2019) adopts bidirectional masked modeling for sequences.
@@
Our reproduced and controlled variants. To isolate architectural effects under identical capacity, tokenizer, and training budget, we implement three LLM-based variants on the same backbone as IDIOMoE: (i) ID Transformer (item tokens only); (ii) Item-ID LLM + text-derived bias (ID embeddings augmented with text features); and (iii) Item-LLM (vocabulary expansion with explicit item text but no MoE). We also reproduce strong non-LLM and hybrid sequential baselines, including SASRec (Kang & McAuley, 2018) and HSTU (Zhai et al., 2024). Unless stated otherwise, all LLM-based baselines are matched to IDIOMoE in active parameter count and trained with the same token budget, optimizer, sequence length, and schedules.
@@ B.6.1 Proprietary Results
B.6.2 Attention analysis on text-only prompts
@@
Figures 6 and 7 show that the MoE model and the pretrained backbone exhibit near-identical attention patterns on text-only inputs across all layers. Layer-wise previous-token bias, first-token emphasis, and attention entropy overlap almost perfectly, and early/mid/late distance profiles coincide within visual resolution.
@@
B.6.3 Efficiency Results
Table 10 summarizes end-to-end efficiency for three sequence lengths. Two trends stand out:

### v2

@@ A.3 Generative and Hybrid Recommender Models
Generative recommenders recast recommendation as a sequence generation task (Yang et al., 2025), unifying retrieval and ranking in one model. HSTU (Zhai et al., 2024) employs a Transformer-based transducer, scaling up to 1.5T parameters and achieving large offline and online gains, while demonstrating NLP-like scaling laws for recommendation. TIGER (Rajput et al., 2023) compresses item vocabularies via multi-code vector quantization. OneRec (Deng et al., 2025b) unifies retrieval and ranking in an encoder–decoder Transformer with sparse Mixture-of-Experts (MoE) Shazeer et al. (2017); Fedus et al. (2022); Ma et al. (2018); Tang et al. (2020); Xu et al. (2024); Zhang et al. (2024); Wang et al. (2024b) for capacity scaling and adds Iterative Preference Optimization for alignment. These approaches offer novelty, explainability, and unified modeling, but require heavy compute and careful fine-tuning strategies to retain collaborative memory.
@@
A.4 Multimodal MoE LLMs
More recently, MoE has been integrated directly into multimodal large language models (MLLMs) and large vision–language models (LVLMs) Bao et al. (2022); Shen et al. (2023); Diao et al. (2025); Deng et al. (2025a); Xiong et al. (2025); Wang et al. (2025b); Xiong et al. (2026); Ganjdanesh et al. (2025); Shirkavand et al. (2025b). MoE-LLaVA (Lin et al., 2024a) introduces a sparse MoE backbone for LLaVA (Liu et al., 2023a)-style LVLMs and proposes a three-stage MoE-tuning strategy that first builds a strong dense LVLM and then converts its feed-forward blocks into experts. The resulting MoE-LLaVA model achieves performance comparable to or better than substantially larger dense LLaVA variants, while activating only a fraction of the parameters per token and reducing visual hallucinations Rawal et al. (2025).
Uni-MoE (Li et al., 2025) targets unified multimodal LLMs that support a broad set of modalities and tasks, applying MoE layers to scale capacity while maintaining a single generalist model. Addressing task interference in instruction-tuned MLLMs, MoME (Mixture of Multimodal Experts)  (Xu et al., 2024) decomposes the architecture into a Mixture of Vision Experts (MoVE) and a Mixture of Language Experts (MoLE). MoVE aggregates features from multiple vision encoders via an adaptive deformable transformation and an instruction-conditioned router, while MoLE inserts sparsely gated adapter experts into LLM layers.
@@ B.1 Baselines
Early sequential modeling. GRU4Rec (Hidasi et al., 2015) pioneers GRU-based session modeling; SASRec (Kang and McAuley, 2018) introduces unidirectional self-attention; BERT4Rec (Sun et al., 2019) adopts bidirectional masked modeling for sequences.
@@
Our reproduced and controlled variants. To isolate architectural effects under identical capacity, tokenizer, and training budget, we implement three LLM-based variants on the same backbone as IDIOMoE: (i) ID Transformer (item tokens only); (ii) Item-ID LLM + text-derived bias (ID embeddings augmented with text features); and (iii) Item-LLM (vocabulary expansion with explicit item text but no MoE). We also reproduce strong non-LLM and hybrid sequential baselines, including SASRec (Kang and McAuley, 2018) and HSTU (Zhai et al., 2024). Unless stated otherwise, all LLM-based baselines are matched to IDIOMoE in active parameter count and trained with the same token budget, optimizer, sequence length, and schedules.
@@ B.6.1 Proprietary Results
We also conduct an additional experiement on our industrial dataset to study the effect of scaling the model using the Qwen 2.5 Qwen et al. (2025) family (0.5B, 1.5B, 3B, 7B). Figure 6 shows the results. We see that recommendation quality improves with LLM size given enough training data, and the gains of IDIOMoE over Item-LLM as the main baseline are persistent across all model scales considered.
Figure 6: Relative performance on the industrial dataset with Qwen 2.5 backbones of different sizes. All values are reported as relative improvements (%) over the 0.5B baseline Item-LLM.
B.6.2 Semantic IDs
Table 10: Performance with Semantic IDs on three Amazon datasets.
Arts Games Instruments
Method HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
Item-LLM 0.0946 0.0658 0.0823 0.0481 0.0826 0.0622
IDIOMoE 0.1018 0.0730 0.0880 0.0492 0.0917 0.0686
IDIOMoE is fully compatible with semantic ID schemes for handling new items. We conduct an experiments where we replace raw item IDs with semantic IDs from MQL4GRec (Zhai et al., 2025), showing that IDIOMoE ’s gains persist in this setting (Table 10).
We also conducted a cold-start experiment following the steps described in the section 4.3 of TIGER (Rajput et al., 2023), where we remove 5% of test items from the training data and report the test performance overall and over the unseen set items. We set the ratio of unseen items to seen items in the top-k items $\epsilon=0.1$. Table 11 shows the results, demonstrating that our method extends naturally to standard cold-start mechanisms.
Table 11: Cold-start evaluation following TIGER: 5% of test items are removed from training. We report overall test metrics (All) and metrics restricted to unseen items (Unseen) for three datasets.
Arts Games Instruments
All Unseen All Unseen All Unseen
Method HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
Backbone 0.0808 0.0618 0.0569 0.0395 0.0849 0.0534 0.0478 0.0332 0.0642 0.0433 0.0394 0.0249
IDIOMoE 0.0892 0.0643 0.0547 0.0416 0.0941 0.0572 0.0541 0.0422 0.0877 0.0579 0.0580 0.0313
These results demonstrate that our method extends naturally to standard cold-start mechanisms and is compatible with semantic-ID-based handling of new items.
B.6.3 Attention analysis on text-only prompts
@@
Figures 7 and 8 show that the MoE model and the pretrained backbone exhibit near-identical attention patterns on text-only inputs across all layers. Layer-wise previous-token bias, first-token emphasis, and attention entropy overlap almost perfectly, and early/mid/late distance profiles coincide within visual resolution.
@@
B.6.4 Efficiency Results
Our model is evaluated with standard batched inference. It is not restricted to processing a single query at a time. Just like a conventional LLM, IDIOMoE supports multi-query batches with appropriate padding and attention masking, and all of our reported results use evaluation batch sizes larger than 1. Table 12 reports latency and throughput values for both batched training and inference on three sequence lengths, showing that IDIOMoE achieves comparable performance values to the underlying backbone model at various sequence lengths. The MoE modification only changes the FFN sublayers (attention remains shared), so the per-token compute remains similar, and there is no additional online adaptation step at serving time beyond a single forward pass. From Table 12 two trends stand out: